OpenStudy (anonymous):

Jackson takes 6 hours to clean the garage alone,but can clean it in 1.5 hours if he works with Mary.how long does it take Mary to clean the garage if she works alone

OpenStudy (md152727):

6/x=1.5

OpenStudy (md152727):

x=4, but I am not positive @AlexandervonHumboldt2

OpenStudy (md152727):

I think teh answer is 4

OpenStudy (md152727):

is there multiple choices?

OpenStudy (anonymous):

yes

OpenStudy (md152727):

can you post them

OpenStudy (anonymous):

\[\ \frac{ 3 }{ 4 } or 2 \frac{ 9 }{ 7} \] 2 is a separate answer